OpenStudy (anonymous):

Find the six roots to the equation ... Find the six roots to the equation ... @Mathematics

OpenStudy (anonymous):

\[z^{6}=-i\]

OpenStudy (phi):

Write -i in polar form exp(i3pi/2)^(1/6) all the answers like equi-spaced on the unit circle

myininaya (myininaya):

i'm sorry zarkon can i borrow you

myininaya (myininaya):

when you get done here of course http://openstudy.com/?F626518757369D0BHIS=_#/updates/4ec0328be4b0caac70006666

OpenStudy (anonymous):

I converted -i to polar form: \[1(\cos \frac{3\pi}{2} + isin \frac{3\pi}{2})\] Now I'm doing: \[1(\cos \frac{\frac{3\pi}{2}+2\pi k}{6} + isin \frac{\frac{3\pi}{2}+2\pi k}{6}\] Where k ranges from 0 to n-1 ( 5 in this case) , am I doing it right?

OpenStudy (phi):

That is still rectangular form. polar form is \[\cos \theta + i \sin \theta =e^{i \theta}\] but you do get all the values doing your way.

OpenStudy (anonymous):

alright, thank you

OpenStudy (phi):

 very rough drawing of the roots starting at 3pi/12 (pi/4), incrementing by 4pi/12 (pi/3)

OpenStudy (anonymous):

pi/3?

OpenStudy (phi):

that's what you get from 2pi k/6 for k=0 to 5

OpenStudy (anonymous):

If the first root is \[1(\cos \frac{\frac{3\pi}{2}+2\pi (0)}{6} + isin \frac{\frac{3\pi}{2}+2\pi (0)}{6}\] That gives cos (pi/4).. i sin(pi/4) So the first root is \[\frac{\sqrt{2}}{2} + \frac{\sqrt{2}}{2}i\] The next one would be \[1(\cos \frac{\frac{3\pi}{2}+2\pi (1)}{6} + isin \frac{\frac{3\pi}{2}+2\pi (1)}{6}\] Wouldn't that be cos (7pi/12) + i sin (7pi/12) and that would be the next root?

OpenStudy (phi):

Yes, or equivalently, e^(i 7pi/12) The angle is pi/4 + pi/3, or 3pi/12 + 4pi/12.

OpenStudy (anonymous):

oh, right I get it now